Janas is a $100 million firm that is currently all equity financed. The firm increases its size from $100 million to $250 million, with all expansion funds coming in the form of new debt. At the new size, which of the following comes closest to the percentage fall in assets before all the equity is wiped out?

0%

25%

40%

60%

100%
Answer #1

%age fall in asset before equity wipe out

=(equity/asset)*100 = (100/250)*100 = 40% fall